The data represent the number of driving fatalities for a certain area by age for male and female drivers.

Male

Female

under 16

114

102

16 dash 20

4198

2139

21 dash 34

14 comma 809

4339

35 dash 54

12 comma 910

5252

55 dash 69

4103

2476

70 and over

2776

1629

​(a) What is the probability that a randomly selected driver fatality who was

female

was

55 to 69

years​ old?
Answer #1

Total female = 102 + 2139 + 4339 + 5252 + 2476 + 1629 = 15937

a) P(55 to 69 years old | female) = 2476 / 15937 = 0.1554